OpenStudy (anonymous):

help no idea how to solve it what is the tenth term of the geometric sequence that has a common ratio of '1/3' and 36 as its fifth term

OpenStudy (anonymous):

A.4/27 B.27/4 C.4/81 D.1/36

OpenStudy (jdoe0001):

\(\huge a_n = a_1(r)^{n-1}\) \(a_1\) = 1st term r = ratio n = the \(n^{th}\) term

OpenStudy (anonymous):

so when you find the n term do you subtract it by 1

OpenStudy (jdoe0001):

so we know that the 5th term is 36, we know that the ratio is 1/3 so what's the first term anyway? well the 5th term equals 36 with a ratio of 1/3, thus $$\bf a_n = a_1(r)^{n-1}\\ a_5 = a_1\left(\cfrac{1}{3}\right)^{5-1}\\ 36 = a_1\left(\cfrac{1}{3}\right)^{4} \implies 36 = a_1\left(\cfrac{1^4}{3^4}\right)\\ \cfrac{ \cfrac{36}{1}}{ \left(\cfrac{1^4}{3^4}\right)} = a_1 $$

OpenStudy (anonymous):

what do you do in the last part

OpenStudy (jdoe0001):

\(\bf \cfrac{ \frac{36}{1}}{ \left(\frac{1^4}{3^4}\right)} = a_1 \implies \cfrac{36}{1} \times \cfrac{81}{1}\)

OpenStudy (jdoe0001):

so now we know that the 1st term, \(a_1 = 36 \times 81 = 2916\)

OpenStudy (jdoe0001):

so you want the \(\bf n^{th}\) term in this case is 10 recall => \(\huge a_n = a_1(r)^{n-1}\)

OpenStudy (jdoe0001):

so, what does that give you?

OpenStudy (anonymous):

im still lost

OpenStudy (jdoe0001):

ok, what confused you? :)

OpenStudy (anonymous):

the formula

OpenStudy (jdoe0001):

ok, what confuses you in the formula?

OpenStudy (anonymous):

like do you multiply everything then subtract the n and 1

OpenStudy (jdoe0001):

hmm, the n-1 is the exponent say if your \(\bf n^{th}\) term is say 23 then n = 23, thus n- 1 => 23 -1

OpenStudy (jdoe0001):

and \(\bf \large a_n = a_{23}\)

OpenStudy (jdoe0001):

n just stands as a generic variable in the notation for the \(\bf n^{th}\) term whichever it happens to be, it can be 3, or 5 or 17, or 1,000,000 in this case it's just 10

OpenStudy (jdoe0001):

still confused?